TWO MODEL HAMILTONIANS FOR TORSION-INVERSION TUNNELING IN THE CH-STRETCH VI-BRATIONALLY EXCITED STATES OF METHYLAMINE

\underline{MAHESH B DAWADI}, AND DAVID S PERRY, Department of Chemistry, The University of Akron, OH 44325-3601.

In methylamine (CH_{3}NH_{2}), there are six equivalent mimina that are connected by torsion and inversion tunneling. In the G_{1}_{2} molecular symmetric group, there are four species, A = {A_{1}, A_{2}}, B = {B_{1}, B_{2}}, E_{1} and E_{2} that combine with distinct nuclear states. The ground vibrational state of CH_{3}NH_{2} is split by torsion and inversion tunneling into a multiplet pattern of four distinct energy levels^{a}. The experimental tunneling pattern for CH_{3}NH_{2} in the ν_{1}_{1} asymmetric CH-stretch fundamental has been previously reported at this meeting. In the experimental pattern, the degenerate species (E_{1} and E_{2}) are at the top and bottom of the multiplet and the non-degenerate species (B and A) are between them. In this work, we present two models for the torsion-inversion tunneling behavior in the CH-stretch excited states. Each model includes the lowest order torsion-inversion-vibration interactions available in the context of the model. The first model, which extends Hougen’s treatment of methanol,^{b} couples the two vibrational angular momentum components of the asymmetric CH-stretches to the large-amplitude motion to yield predicted tunneling patterns for the ν_{2} and ν_{1}_{1} fundamentals. This model gives similar patterns for ν_{2} and ν_{1}_{1}, in which E_{1} and E_{2} are in the middle of the multiplet and the non-degenerate species are at the top and bottom. The second model, which follows conceptually Wang and Perry’s local mode treatment of methanol,^{c} couples the three local CH-stretches to each other and to the large-amplitude motion to yield the tunneling patterns for the ν_{2}, ν_{3} and ν_{1}_{1} fundamentals. For this model, we found that, for ν_{2} and ν_{1}_{1}, both E_{1} and E_{2} are at the bottom of the multiplet, in contrast to ν_{3} and the ground state where they are at the top. The fact that neither model reproduces the observed tunneling pattern for ν_{1}_{1}, suggests that additional isolated perturbations or systematic interactions are present in the experimental spectra.

aV.V. Ilyshin et al.J. Mol. Spectrosc. \underline{251}(56-63), 2008.

bJ.T. Hougen J. Mol. Spectrosc. \underline{207}(60-65), 2001.

cX. Wang and D.S. Perry J. Chem. Phys. \underline{109}(10795-10805), 1998.